If your uncle is 2 years older than 3 times your age, how old is he?
Found 2 solutions by fractalier, addingup:
Answer by fractalier(6550)

All depends how old you are...take your age, multiply it times three and then add two...

Answer by addingup(3677)

The way you wrote it the problem has no meaningful solution.
You say: Uncle= 2+3*your age. Look what happens, you can make it just about anything you want:
---------------------------------
You:10, Uncle: 2+3(10)= 32
You:11, Uncle: 2+3(11)= 35
You:12, Uncle: 2+3(12)= 38
etcetera
---------------
You are missing a key piece of information, an equality to balance the equation. that's what equation means, it comes from equal. Check your problem and resubmit it.
J